2x+3(x-5)=15-(x-12)
Answer by nyc_function(2741)

2x+3(x-5)=15-(x-12)
2x + 3x - 15 = 15 - x + 12
5x - 15 = -x + 27
5x + x = 15 + 27
6x = 42
x = 42/6
x = 7
Done!